On January 1, Year 1, Price Co. issued $144,000 of five-year, 6 percent bonds at 95. Interest is payable annually on December 31. The discount is amortized using the straight-line method. Required Prepare the journal entries to record the bond transactions for Year 1 and Year 2. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

A) Record the entry for issuance of bonds.

B) Record the entry for recognizing interest expense on Dec. 31, Year 1.

C) Record the entry for recognizing interest expense on Dec. 32, Year 2.
Answer #1

Face Value of bonds = $ 144000

issue price of the bond = $ 95

no of bonds issued = 144000 / 100

= 1440 bonds

proceeds from issue of bonds = 1440 * 95

= $ 136800

discount on bonds = Face Vaalue - proceeds From issue

= 144000 - 136800

= $ 7200

Straight line Amortization per year

= Discount / Term of the bonds

= 7200 / 5

= $ 1440 per year

interest payment for Annual period = 144000 * 6%

= $ 8640

Date Accounts Name Debit Credit
A Bond Issue journal entry  
1-Jan Cash 136800
Discount on bonds 7200
      Bonds Payable 144000
B
year - 1 Interest Expense ( 8640+1440) 10080
Discount on bonds 1440
         Cash    8640
year 2   
C Interest Expense ( 8640+1440) 10080
Discount on bonds 1440
         Cash    8640